OpenStudy (anonymous):

Which of the following points lies on the circle whose center is at the origin and whose radius is 10? (6, -8)

OpenStudy (amistre64):

this is just the pythag thrm at work such that: x^2 + y^2 = r^2

OpenStudy (amistre64):

you can test the options by squaring the point values, and adding them together, if they equate to teh square of the radius ... its good

OpenStudy (anonymous):

so did i got it right?

OpenStudy (amistre64):

lets chk, what is: 6^2 (-8)^2 10^2

OpenStudy (anonymous):

36 64 100..

OpenStudy (amistre64):

good, and since 36+64 = 100; its a perfect fit